A person travels halff the total dist. at 30 km/hr and the next half at 20 km/hr. find the average velocity (SOLVE USING FORMULA AND SHOW THE DERIVATION OF FORMULA>)

Grade:10

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Let x be the total distance.

time taken t1 for 1st part=x/2*30=x/60

t2=x/40

total time taken=x/60+x/40= 2x+3x/120=5x/120

total distance=x

average velocity=x*120/5x=24 km/hr

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi